# DEEP-ICL: Definition-Enriched Experts for Language Model In-Context Learning

### v1

@@ Abstract
It has long been assumed that the sheer number of parameters in large language models (LLMs) drives in-context learning (ICL) capabilities, enabling remarkable performance improvements by leveraging task-specific demonstrations. Challenging this hypothesis, we introduce DEEP-ICL, a novel task Definition Enriched ExPert Ensembling methodology for ICL. DEEP-ICL explicitly extracts task definitions from given demonstrations and generates responses through learning task-specific examples. We argue that improvement from ICL does not directly rely on model size, but essentially stems from understanding task definitions and task-guided learning. Inspired by this, DEEP-ICL combines two 3B models with distinct roles (one for concluding task definitions and the other for learning task demonstrations) and achieves comparable performance to LLaMA2-13B. Furthermore, our framework outperforms conventional ICL by overcoming pretraining sequence length limitations, by supporting unlimited demonstrations. We contend that DEEP-ICL presents a novel alternative for achieving efficient few-shot learning, extending beyond the conventional ICL.
DEEP-ICL: Definition-Enriched Experts for Language Model
In-Context Learning
thanks: Equal contributionXingwei Qu2,311footnotemark: 1Yiming Liang4,5,10 Yucheng Wang1 Tianyu Zheng1,9 Tommy Yue7Lei Ma8,10Stephen W. Huang11Jiajun Zhang4,5Wenhu Chen6 thanks: Corresponding author.Chenghua Lin322footnotemark: 2Jie Fu1,211footnotemark: 122footnotemark: 2Ge Zhang1,2,6 1M-A-P 2HKUST  3 University of Manchester 4Institute of Automation, CAS 5University of Chinese Academy of Sciences 6University of Waterloo 7Ohio State University 8Peking University 9Beijing University of Posts and Telecommunications 10BAAI 11Harmony.ai ge.zhang@uwaterloo.ca, jiefu@ust.hk, chenghua.lin@manchester.ac.uk
@@ 1 Introduction
Figure 1: Pipeline of DEEP-ICL. The preparation stage focuses on generating a pool of task-based experts, each trained using task-specific data. ICL firstly generates task definitions from given demonstrations using a task definition generator, which can be a custom generator, GPT, or human. Secondly, we utilize the task definitions to retrieve the task-based expert pool and match the top three experts that are most similar. Next, the expert ensemble step initializes the fusion expert by performing a weighted average of the matched experts’ weights. Finally, the data format is reconstructed for training the fusion expert and performing inference on the last input field.
In recent years, large language models (LLMs) have showcased remarkable capabilities, particularly for in-context learning (ICL). This method allows LLMs to learn how to execute specific tasks by analyzing a few labeled examples provided in their input context, without any updates to their parameters. This breakthrough enables LLMs to serve as versatile, general-purpose models capable of tackling a wide range of tasks with minimal guidance, as highlighted by Brown et al. (2020).
@@
In this paper, we propose a novel ICL-style ensembling approach through the lens of few-shot learning, called task Definition Enriched ExPert-ICL (DEEP-ICL). Inspired by Pan et al. (2023), our approach incorporates two 3B models: one focuses on task definition determination, and the other specializes in learning from task-specific demonstrations. Since few-shot learning usually targets specific tasks, it is difficult to generalize to other unseen tasks. To overcome this problem, we create a dynamic expert pool filled with prior knowledge during the demonstration learning phase. In other words, the experts in the pool can be selected according to the task definition derived in the task definition generation phase. Furthermore, this dynamic pool utilizes LoRA (Hu et al., 2021) for selecting the top experts based on task requirements. Our experiments show that even with continual training on just five samples, our ensembling methods outperform both traditional ICL and non-ensembling approaches. This setup supports lifelong learning by surpassing the fixed limitations of LLMs, significantly improving performance.
We evaluate DEEP-ICL on 117 subtasks in the official Super Natrual Instruction (SuperNI) Wang et al. (2022b) test dataset split Wang et al. (2022b). Our experiments, comparing the performance to LLaMA2 models with 7B and 13B parameters Touvron et al. (2023), and Falcon 7B Almazrouei et al. (2023), show that our models significantly outperform these two 7B level models and exhibit performance that is marginally below the LLaMA2-13B model, indicating our approach is competitive within a narrow margin. This result demonstrates that our solution is effective in addressing the limitations of ICL and few-shot learning techniques.
@@ 2.3 Task Definition Extraction
Our methodology, DEEP-ICL, aligns with the prevailing emphasis on instructions for task definition in LLMs. Unlike traditional methods focusing on autonomous instruction generation or refinement, DEEP-ICL integrates task definition extraction with expert ensembling. This approach ensures optimal model performance by accurately interpreting task definitions and distinguishing DEEP-ICL as a more efficient and adaptive solution in the LLM landscape.
@@ 3 Methods
DEEP-ICL represents a model ensembling method for generalizable and continual few-shot learning. This section delves into the details about how DEEP-ICL performs few-shot learning.
@@ 4 Experimental Setup
4.1 Evaluating the Performance of DEEP-ICL
@@
Table 1: The performance of DEEP-ICL across different task categories. K represents the number of samples. Due to space constraints, we select scores from both the 5-sample and the 500-sample settings for result demonstration. For more details, please refer to Appendix Table 6.
We conduct experiments based on SuperNI, a dataset comprised of human-authored instructions, which encompasses a variety of tasks, to verify the generalization and performance enhancement of DEEP-ICL. We point out that only the parameters of the task definition extractor, task definition guided retriever, and ensembled expert are activated during the inference pipeline of DEEP-ICL. The estimated activated parameter size of DEEP-ICL is around 6B if we utilize our 3B paradigm. As a result, we select LLaMA2-7B, LLaMA2-13B, and Falcon-7B as our baselines, which the most widely utilized open-source models, demonstrating that DEEP-ICL achieves comparable few-shot performance to LLMs twice its size.
In this section, we detail the SuperNI-based experiment setting. Firstly, we construct an expert pool. For this purpose, we employ the T5 model series (Raffel et al., 2020) as our expert base model. Each T5 expert is specifically trained with LoRA (Hu et al., 2021) for each task from SuperNI official training splits to constitute DEEP-ICL’s prior knowledge. Due to many tasks being smaller in size, we opt for a 9:1 split for training and validation. Had we gone for an 8:2 train-test split, our overall training set would have been insufficient. For each task, adaptors and instructions are extracted from SuperNI. During training, the batch size is set to 40, the learning rate at 3e-4, the input length is 1024, and the LoRA trained parameters are the "q_proj" and k_proj modules in the multi-head attention modules.
@@ 4.2 Ablation Experiment
We conduct two ablation experiments to investigate the roles of the task definition and expert ensemble components in the DEEP-ICL process. We reconstructed the input-output data of demonstrations and created two formats of labeled data to train LoRA in two different ways.
The setting of Deep-ICL without definitions and ensemble, as stated in Table 2, refers to directly using few-shot input-output pairs as labeled data to train randomly initialized experts. In this scenario, the DEEP-ICL stage lacks both task definitions derived from user demonstrations and expert ensembling and degenerates into few-shot learning. We conduct these ablation experiments on settings with 5 samples, 50 samples, 100 samples, and 500 samples, as well as using T5 small, T5 base, T5 large, and T5 3B models, to investigate the impact of task definition in Deep-ICL under different data scales and model sizes.
The ablation experiment setting of Deep-ICL without ensemble firstly retrieves task definition from user demonstrations using our own 3B extractor. Next, we add the task definition into the input field in the user demonstrations and construct input-output pairs as labeled data for training LoRA. Compared to standard DEEP-ICL, DEEP-ICL without ensembling lacks the initialization of LoRA weights through expert ensembling and trains LoRA from scratch. We conduct the same experiments as DEEP-ICL without definitions and ensembling to explore the influence of expert ensembling in DEEP-ICL under different data scales and model sizes.
@@ 4.3 Analysing Task Definition Quality
DEEP-ICL retrieves from the expert pool using the generated task definition and matches the top 3 experts based on similarity. This similarity is computed by using the task definition generated from user input and the task definitions stored in the expert pool. Automatic quality analysis approach counts the number of overlapping matched experts as a measure of the commonality of task definitions from different sources, such as task definitions generated by GPT-3, GPT-3.5, GPT-4, our extractor, and those stored in the expert pool.
@@ 5 Results
Figure 2: Performance comparison of DEEP-ICL across four distinct expert base model sizes and five task definition extractors with 5 demonstration inputs. The expert base models include LLaMA2-7B and LLaMA2-13B, and Falcon7B. Baseline ICL results are denoted by dashed lines. Notably, our paradigm employing a 6B parameters demonstrates superior performance to the LLaMA2-7B ICL and exhibits comparable results to the LLaMA2-13B ICL.
5.1 Performance of DEEP-ICL
@@ 5.2 Ablation Study
Table 2: Ablation results. We conduct ablation experiments on T5_small, T5_base, T5_large, and T5_3B models, using 5 samples, 50 samples, 100 samples, and 500 samples. Olive color values and cyan color values indicate ablations of DEEP-ICL without definitions and ensembling and DEEP-ICL without ensembling, respectively. TDG refers to task definition generator.
Compared to the final DEEP-ICL, DEEP-ICL without definitions and ensembling exhibits an average performance decrease of 44.71% across various data scales and model sizes. This underscores the significant role of task definitions in DEEP-ICL, particularly within demonstrations of the same task type. In addition, another observation is that DEEP-ICL without definitions and ensembling does not show significant performance improvements as the data scale and model size increases, indicating that without explicit (ours) or implicit (ICL) task definition extraction, few-shot learning does not work. From another perspective, it underscores the crucial role of task definitions for this paradigm, both in ICL and in our approach.
In the ablation experiment setting of DEEP-ICL without ensembling, there are four experimental setups ([T5 small, 5 samples], [T5 small, 50 samples], [T5 base, 5 samples], [T5 large, 5 samples]) that perform more than 10% worse than the final DEEP-ICL. This indicates that in scenarios with small models or limited data, experts ensemble in DEEP-ICL facilitates rapid bootstrapping of inference for related tasks.
As the number of samples increases and the model size grows, the performance gap between final DEEP-ICL and DEEP-ICL without ensembling gradually diminishes. This suggests that the initialization of DEEP-ICL by using expert ensembling does not negatively impact the effectiveness of model finetuning in the future.
@@ 6 Conclusion
In this paper, we present DEEP-ICL, a new approach that addresses the limitations of existing ICL frameworks related to fixed input lengths and limited learning capabilities. DEEP-ICL utilizes an expert pool to provide a strong knowledge base and improve learning from input demonstrations. We divide the ICL process into task definition extraction and task processing. Our findings show that accurately extracting task definitions is vital for ICL success. During task processing, we employ a novel initialization technique based on task definitions, leveraging insights from the expert pool, resulting in significantly enhanced performance and advancing the ICL process.
@@ Appendix C The Table of Model Performance
Table 5 shows the performance of DEEP-ICL under different experimental settings, such as task definition generator, LoRA base model, and number of ICL examples. The human setting is to directly use the task definitions in the dataset.
@@
Table 6 displays the performance of DEEP-ICL across different task categories in detail.
@@
Table 5: Performance of DEEP-ICL under different experimental settings, such as Task Definition Generator, LoRA base model, and number of ICL examples. The Human setting is to directly use the task definitions in the dataset. The No Definition setting is to directly train adapter soup without using task definition. The best results in each section are Bold, the second-best results are underlined.
(a) GPT3 & GPT-3.5
(b) GPT3 & GPT4
(c) GPT3 & Human
(d) GPT3 & Ours
(e) GPT-3.5 & GPT4
(f) GPT-3.5 & Human
(g) GPT-3.5 & Ours
(h) GPT4 & Human
(i) GPT4 & Ours
@@
Table 6: The performance of DEEP-ICL across different task categories.

### v3

# TEGEE: Task dEfinition Guided Expert Ensembling for Generalizable and Few-shot Learning

@@ Abstract
Large Language Models (LLMs) exhibit the ability to perform in-context learning (ICL), where they acquire new tasks directly from examples provided in demonstrations. This process is thought to operate through an implicit task selection mechanism that involves extracting and processing task definitions from these demonstrations. However, critical questions remain: Which is more essential—task extraction or definition? And how can these capabilities be further improved? To address these questions, we propose TEGEE (Task Definition Guided Expert Ensembling), a method that explicitly extracts task definitions and generates responses based on specific tasks. Our framework employs a dual 3B model approach, with each model assigned a distinct role: one focuses on task definition extraction, while the other handles learning from demonstrations. This modular approach supports the hypothesis that extracting task definitions is more vital than processing the task itself. Empirical evaluations show that TEGEE performs comparably to the larger LLaMA2-13B model. By leveraging a modular design, our approach extends traditional ICL from few-shot to many-shot learning, supporting an unlimited number of demonstrations and enhancing continual learning capabilities.
TEGEE: Task dEfinition Guided Expert Ensembling for Generalizable and Few-shot Learning
thanks: Equal contributionXingwei Qu2,311footnotemark: 1Yiming Liang4,5,10 Yucheng Wang1 Tianyu Zheng1,9 Tommy Yue7Xingyuan Bu1Lei Ma8,10Stephen W. Huang11Jiajun Zhang4,5 Yinan Shi1thanks: Corresponding author.Chenghua Lin322footnotemark: 2Jie Fu211footnotemark: 122footnotemark: 2Ge Zhang1,2,6 1M-A-P 2HKUST  3 University of Manchester 4Institute of Automation, CAS 5University of Chinese Academy of Sciences 6University of Waterloo 7Ohio State University 8Peking University 9Beijing University of Posts and Telecommunications 10BAAI 11Harmony.ai ge.zhang@uwaterloo.ca, jiefu@ust.hk, chenghua.lin@manchester.ac.uk
@@ 1 Introduction
Figure 1: Pipeline of TEGEE. The preparation stage focuses on generating a pool of task-based experts, each trained using task-specific data. ICL firstly generates task definitions from given demonstrations using a task definition generator, which can be a custom generator, GPT, or human. Secondly, we utilize the task definitions to retrieve the task-based expert pool and match the top three experts that are most similar. Next, the expert ensemble step initializes the fusion expert by performing a weighted average of the matched experts’ weights. Finally, the data format is reconstructed for training the fusion expert and performing inference on the last input field.
In recent years, large language models (LLMs) have showcased remarkable capabilities, particularly for in-context learning (ICL). This method allows LLMs to learn how to execute specific tasks by analyzing a few labeled examples provided in their input context, without any updates to their parameters. This breakthrough enables LLMs to serve as versatile, general-purpose models capable of tackling a wide range of tasks with minimal guidance, as highlighted by Brown et al. (2020). However, there is still no consensus on when or how ICL works.
@@
In this paper, we propose a novel ICL-style ensembling approach through the lens of few-shot learning, called Task dEfinition Guided Expert Ensembling (TEGEE). Inspired by Pan et al. (2023), our approach incorporates two 3B models: one focuses on task definition determination, and the other specializes in learning from task-specific demonstrations. Since few-shot learning usually targets specific tasks, it is difficult to generalize to other unseen tasks. To overcome this problem, we create a dynamic expert pool filled with prior knowledge during the demonstration learning phase. In other words, the experts in the pool can be selected according to the task definition derived in the task definition generation phase. Furthermore, this dynamic pool utilizes LoRA (Hu et al., 2021) for selecting the top experts based on task requirements. Our experiments show that even with continual training on just five samples, our ensembling methods outperform both traditional ICL and non-ensembling approaches. This setup supports lifelong learning by surpassing the fixed limitations of LLMs, significantly improving performance.
We evaluate TEGEE on 117 subtasks in the official Super Natrual Instruction (SuperNI) Wang et al. (2022b) test dataset split Wang et al. (2022b). Our experiments, comparing the performance to LLaMA2 models with 7B and 13B parameters Touvron et al. (2023), and Falcon 7B Almazrouei et al. (2023), show that our models significantly outperform these two 7B level models and exhibit performance that is marginally below the LLaMA2-13B model, indicating our approach is competitive within a narrow margin. This result demonstrates that our solution is effective in addressing the limitations of ICL and few-shot learning techniques.
@@ 2.3 Task Definition Extraction
Our methodology, TEGEE, aligns with the prevailing emphasis on instructions for task definition in LLMs. Unlike traditional methods focusing on autonomous instruction generation or refinement, TEGEE integrates task definition extraction with expert ensembling. This approach ensures optimal model performance by accurately interpreting task definitions and distinguishing TEGEE as a more efficient and adaptive solution in the LLM landscape.
@@ 3 Methods
TEGEE represents a model ensembling method for generalizable and continual few-shot learning. This section delves into the details about how TEGEE performs few-shot learning.
@@ 4 Experimental Setup
4.1 Evaluating the Performance of TEGEE
@@
Table 1: The performance of TEGEE across different task categories. K represents the number of samples. Due to space constraints, we select scores from both the 5-sample and the 500-sample settings for result demonstration. For more details, please refer to Appendix Table 6.
We conduct experiments based on SuperNI, a dataset comprised of human-authored instructions, which encompasses a variety of tasks, to verify the generalization and performance enhancement of TEGEE. We point out that only the parameters of the task definition extractor, task definition guided retriever, and ensembled expert are activated during the inference pipeline of TEGEE. The estimated activated parameter size of TEGEE is around 6B if we utilize our 3B paradigm. As a result, we select LLaMA2-7B, LLaMA2-13B, and Falcon-7B as our baselines, which the most widely utilized open-source models, demonstrating that TEGEE achieves comparable few-shot performance to LLMs twice its size.
In this section, we detail the SuperNI-based experiment setting. Firstly, we construct an expert pool. For this purpose, we employ the T5 model series (Raffel et al., 2020) as our expert base model. Each T5 expert is specifically trained with LoRA (Hu et al., 2021) for each task from SuperNI official training splits to constitute TEGEE’s prior knowledge. Due to many tasks being smaller in size, we opt for a 9:1 split for training and validation. Had we gone for an 8:2 train-test split, our overall training set would have been insufficient. For each task, adaptors and instructions are extracted from SuperNI. During training, the batch size is set to 40, the learning rate at 3e-4, the input length is 1024, and the LoRA trained parameters are the "q_proj" and k_proj modules in the multi-head attention modules.
@@ 4.2 Ablation Experiment
We conduct two ablation experiments to investigate the roles of the task definition and expert ensemble components in the TEGEE process. We reconstructed the input-output data of demonstrations and created two formats of labeled data to train LoRA in two different ways.
The setting of TEGEE without definitions and ensemble, as stated in Table 2, refers to directly using few-shot input-output pairs as labeled data to train randomly initialized experts. In this scenario, the TEGEE stage lacks both task definitions derived from user demonstrations and expert ensembling and degenerates into few-shot learning. We conduct these ablation experiments on settings with 5 samples, 50 samples, 100 samples, and 500 samples, as well as using T5 small, T5 base, T5 large, and T5 3B models, to investigate the impact of task definition in TEGEE under different data scales and model sizes.
The ablation experiment setting of TEGEE without ensemble firstly retrieves task definition from user demonstrations using our own 3B extractor. Next, we add the task definition into the input field in the user demonstrations and construct input-output pairs as labeled data for training LoRA. Compared to standard TEGEE, TEGEE without ensembling lacks the initialization of LoRA weights through expert ensembling and trains LoRA from scratch. We conduct the same experiments as TEGEE without definitions and ensembling to explore the influence of expert ensembling in TEGEE under different data scales and model sizes.
@@ 4.3 Analysing Task Definition Quality
TEGEE retrieves from the expert pool using the generated task definition and matches the top 3 experts based on similarity. This similarity is computed by using the task definition generated from user input and the task definitions stored in the expert pool. Automatic quality analysis approach counts the number of overlapping matched experts as a measure of the commonality of task definitions from different sources, such as task definitions generated by GPT-3, GPT-3.5, GPT-4, our extractor, and those stored in the expert pool.
@@ 5 Results
Figure 2: Performance comparison of TEGEE across four distinct expert base model sizes and five task definition extractors with 5 demonstration inputs. The expert base models include LLaMA2-7B and LLaMA2-13B, and Falcon7B. Baseline ICL results are denoted by dashed lines. Notably, our paradigm employing a 6B parameters demonstrates superior performance to the LLaMA2-7B ICL and exhibits comparable results to the LLaMA2-13B ICL.
5.1 Performance of TEGEE
@@ 5.2 Ablation Study
Table 2: Ablation results. We conduct ablation experiments on T5_small, T5_base, T5_large, and T5_3B models, using 5 samples, 50 samples, 100 samples, and 500 samples. Olive color values and cyan color values indicate ablations of TEGEE without definitions and ensembling and TEGEE without ensembling, respectively. TDG refers to task definition generator.
Compared to the final TEGEE, TEGEE without definitions and ensembling exhibits an average performance decrease of 44.71% across various data scales and model sizes. This underscores the significant role of task definitions in TEGEE, particularly within demonstrations of the same task type. In addition, another observation is that TEGEE without definitions and ensembling does not show significant performance improvements as the data scale and model size increases, indicating that without explicit (ours) or implicit (ICL) task definition extraction, few-shot learning does not work. From another perspective, it underscores the crucial role of task definitions for this paradigm, both in ICL and in our approach.
In the ablation experiment setting of TEGEE without ensembling, there are four experimental setups ([T5 small, 5 samples], [T5 small, 50 samples], [T5 base, 5 samples], [T5 large, 5 samples]) that perform more than 10% worse than the final TEGEE. This indicates that in scenarios with small models or limited data, experts ensemble in TEGEE facilitates rapid bootstrapping of inference for related tasks.
As the number of samples increases and the model size grows, the performance gap between final TEGEE and TEGEE without ensembling gradually diminishes. This suggests that the initialization of TEGEE by using expert ensembling does not negatively impact the effectiveness of model finetuning in the future.
@@ 6 Conclusion
In this paper, we present TEGEE, a new approach that addresses the limitations of existing ICL frameworks related to fixed input lengths and limited learning capabilities. TEGEE utilizes an expert pool to provide a strong knowledge base and improve learning from input demonstrations. We divide the ICL process into task definition extraction and task processing. Our findings show that accurately extracting task definitions is vital for ICL success. During task processing, we employ a novel initialization technique based on task definitions, leveraging insights from the expert pool, resulting in significantly enhanced performance and advancing the ICL process.
@@ Appendix C The Table of Model Performance
Table 5 shows the performance of TEGEE under different experimental settings, such as task definition generator, LoRA base model, and number of ICL examples. The human setting is to directly use the task definitions in the dataset.
@@
Table 6 displays the performance of TEGEE across different task categories in detail.
@@
Table 5: Performance of TEGEE under different experimental settings, such as Task Definition Generator, LoRA base model, and number of ICL examples. The Human setting is to directly use the task definitions in the dataset. The No Definition setting is to directly train adapter soup without using task definition. The best results in each section are Bold, the second-best results are underlined.
@@
Table 6: The performance of TEGEE across different task categories.